### v2

@@ Abstract
This work presents a Total Lagrangian (TL) finite element formulation for deformable body dynamics. We employ the TL-FEA framework to simulate the time evolution of collections of bodies whose motion is constrained by kinematic constraints and which mutually interact through contact and friction. These bodies experience large displacements, large deformations, and large rotations. A systematic approach is proposed for classifying and posing kinematic constraints acting between the bodies present in the system. We derive the governing equations for ANCF beam, ANCF shell, and tetrahedral elements, and present hyperelastic material models including St. Venant–Kirchhoff and Mooney–Rivlin formulations with their corresponding internal force contributions and consistent tangent stiffness matrices. A finite-strain Kelvin–Voigt viscous damping model is incorporated in the TL-FEA formulation for numerical stability.
1 Background and Related Work
Large-deformation deformable-body dynamics can be posed naturally within a continuum mechanics setting by describing kinematics through the deformation map $\mathbf{r}(\mathbf{X},t)$ and its associated deformation gradient $\mathbf{F}(\mathbf{X},t)=\partial\mathbf{r}/\partial\mathbf{X}$. In a Total Lagrangian (TL) formulation, all strain measures, stress measures, and constitutive responses are expressed with respect to the reference configuration, leading to a conceptually clean separation between geometry (reference domain), kinematics (through $\mathbf{F}$ and its invariants), and material response (through a strain energy density and/or rate-dependent extensions). This deformation-gradient-centric viewpoint is standard in nonlinear finite elements for solids and structures and provides a universal interface for accommodating hyperelasticity, finite-strain viscous regularization, and multi-field extensions while retaining a consistent linearization and tangent structure Bonet2008nonlinear .
Here, $\mathbf{X}$ denotes generic reference/material coordinates. In element-level derivations, we parameterize the reference domain with local coordinates $\mathbf{u}=[u,v,w]^{T}$, so that locally $\mathbf{X}\equiv\mathbf{u}$.
Within flexible multibody dynamics, the Absolute Nodal Coordinate Formulation (ANCF) introduced by Shabana was a key step toward combining large rigid-body motion with large deformation in a non-incremental finite-element setting Shabana1996 . The defining feature of ANCF is the choice of nodal coordinates: instead of interpolating rotations, ANCF employs absolute position vectors and their spatial gradients as nodal degrees of freedom. This coordinate selection yields a polynomial representation of the position field that remains valid under arbitrarily large rotations and deformations, and it enables a constant (configuration-independent) consistent mass matrix for many element families Shabana1998 . In Shabana’s original presentation and in the standard ANCF literature that followed, the position field is written in the familiar FE form $\mathbf{r}(\mathbf{X},t)=\mathbf{S}(\mathbf{X})\mathbf{e}(t)$, where $\mathbf{S}(\mathbf{X})$ is a block-structured shape-function matrix and $\mathbf{e}(t)$ stacks nodal positions and slopes into a generalized coordinate vector shabanaYakoub2001Part1 ; Shabana2023Overview ; RecueroNegrut2016ANCF .
A large body of work has since expanded ANCF element technology and clarified its modeling tradeoffs. Early developments addressed the construction of beam and cable elements that admit large axial and bending deformation while capturing geometrically nonlinear coupling effects without resorting to rotation interpolation berzeri2000 . Subsequent research introduced shear-deformable and higher-order elements and investigated numerical pathologies such as shear and membrane locking, together with formulation-level remedies that retain ANCF kinematics while improving element accuracy and robustness GarciaVallejo2007LockingFree . Comprehensive surveys document these developments and highlight a recurring theme: although ANCF elements come in many variants (beams, plates/shells, solid-like continua), their kinematics can be consistently viewed as a continuum mapping whose primary computational artifact is the deformation gradient and its derived strain measures Gerstmayr2013 ; Otsuka2022RecentAdvances ; Shabana2023Overview .
[…]

### v3

@@ Abstract
We present a Total Lagrangian finite element framework for finite-deformation multibody dynamics. The framework combines a compact kinematic representation, a deformation-gradient-based formulation, an element-agnostic constitutive interface, and a systematic constraint-construction machinery for coupling deformable bodies through engineering joints. Within this setting, we derive the equations of motion for collections of deformable bodies, and formulate their response in the presence of external loads, frictional contact forces, and constraint reaction forces. The framework accommodates field forces applied pointwise, over surfaces, or throughout volumes, and supports material models of practical interest, including Mooney-Rivlin, Neo-Hookean, and Kelvin-Voigt. A companion paper discusses the GPU-accelerated implementation of the framework outlined herein, and reports on numerical experiments and benchmark results.
keywords:
Total Lagrangian , finite element analysis , multibody dynamics , hyperelasticity , augmented Lagrangian
2020 MSC:
74B20 , 74S05 , 74H55
journal: Computer Methods in Applied Mechanics and Engineering
1 Introduction
This paper presents a Total Lagrangian finite element formulation for constrained multibody dynamics with finite deformations. The focus is on the formulation itself: kinematics, constitutive response, external loading, frictional contact, and bilateral constraints are written in a common notation that is compatible with implicit time integration. A companion paper, submitted simultaneously and available as a preprint [json-ganesh-dan-arxiv-2026feaII], addresses numerical methods, GPU implementation, and benchmark problems built on this formulation.
In the Total Lagrangian setting, all kinematic and stress quantities are referred to the reference configuration. This choice is standard in geometrically nonlinear solid mechanics [1, 2] and leads to a clean separation between reference-domain geometry and deformation-dependent response through the deformation gradient $\mathbf{F}$ and derived strain measures. It also allows derivatives of the interpolation to be evaluated with respect to the reference domain, which is attractive in large-deformation settings; see, e.g., [3] for comparisons against the Updated Lagrangian and corotational alternatives.
Within flexible multibody dynamics, the Absolute Nodal Coordinate Formulation (ANCF) [4, 5] can be viewed as a particular instance of this broader Total Lagrangian perspective. ANCF uses absolute positions and their spatial gradients as nodal coordinates, which yields a polynomial position field valid under large rotations and deformations, produces a constant mass matrix, and uses the same deformation gradient $\mathbf{F}$ as standard continuum finite elements. A related variant by Zhang [6] eliminates the deformation gradient entirely in an explicit, single-body setting; the present work retains $\mathbf{F}$ as the central kinematic quantity, since implicit constrained multibody dynamics requires both the tangent stiffness and the constraint Jacobians that derive from it.
A central modeling choice concerns how the nodal unknowns enter the position field. In much of the finite element and ANCF literature, the interpolation is written in the form $\mathbf{r}=\mathbf{S}(\mathbf{u})\,\mathbf{e}(t)$, where $\mathbf{u}$ denotes the reference coordinate, $\mathbf{S}$ is assembled from shape functions, and $\mathbf{e}$ collects the nodal unknowns [1, 7]. Here, the same interpolation is written in the equivalent compact form $\mathbf{r}=\mathbf{N}(t)\,\mathbf{s}(\mathbf{u})$, where $\mathbf{N}$ stores the nodal unknowns as columns and $\mathbf{s}$ contains the scalar shape functions. In this notation, the deformation gradient takes the form $\mathbf{F}=\mathbf{N}(t)\,\mathbf{H}(\mathbf{u})$, with $\mathbf{H}$ collecting derivatives of the shape functions with respect to the reference coordinates. Earlier contributions in this direction include the edge-matrix form used for linear tetrahedra [8] and the component-free Lagrangian formulation of [9], which avoids Voigt notation but retains the tensor summation form. The functional value of this representation is that it renders the constraint Jacobian derivation element-agnostic: because $\mathbf{F}$, the internal force gradient, and the first variation of the position field all factor through the same $\mathbf{N}$$\mathbf{H}$ split, constraint derivatives across all element types follow from a single chain-rule pattern applied to the point-evaluation operator, without element-specific case analysis.
The treatment of kinematic constraints is a central part of the formulation. In rigid-body dynamics, primitive constraints such as DP1, DP2, CD, and DIST [10] provide a standard route to the construction of engineering joints. For deformable bodies, Betsch and Steinmann [11] developed rotationless constrained dynamics for geometrically exact beams, but their treatment is specific to that element class and does not extend to isoparametric solid or shell elements, nor does it address the Jacobian accumulation and conditioning issues that arise when dot-product and coordinate-difference constraints are mixed. Sugiyama et al. [12] formulated ANCF joint constraints using position-gradient degrees of freedom as orientation surrogates, providing a foundation for deformable-body joints; however, their derivation is cast in the block shape-function notation of ANCF and does not supply the element-level Jacobian accumulation rules, the curvature contribution to the Newton system, or the row-scaling analysis needed for a well-conditioned mixed constraint set. The present work addresses these gaps by applying the primitive-constraint viewpoint directly to isoparametric finite elements in the $\mathbf{N}(t)\,\mathbf{s}(\mathbf{u})$ setting, deriving explicit element-level Jacobian blocks for each primitive, characterizing the DP1 Hessian structure, and providing a conditioning analysis for mixed CD/DP1 joint systems.
On the constitutive side, the formulation uses the first Piola–Kirchhoff stress $\mathbf{P}(\mathbf{F})$ and its material derivative $\partial\mathbf{P}/\partial\mathbf{F}$ as the interface between the constitutive law and the spatial discretization [1, 13, 14]. This separates constitutive modeling from element topology in a natural way. Within this setting, the paper derives nodal internal-force expressions and consistent linearizations for several hyperelastic and viscoelastic material models. In the ANCF literature, García-Vallejo et al. [15] derived elastic force Jacobians for the St. Venant–Kirchhoff model, while subsequent work on nonlinear constitutive models [16, 17] relied on numerical tangent approximations. The present work provides closed-form tangents for the Mooney–Rivlin and Kelvin–Voigt models in the same $\mathbf{P}(\mathbf{F})$ interface, making them directly available for any element topology supported by the framework.
The contributions of this paper are as follows. First, it derives a systematic treatment of deformable-body joint constraints for isoparametric finite elements, including explicit element-level Jacobian accumulation rules for all primitive constraint types, the consistent Newton linearization with the DP1 curvature term, and a conditioning analysis for mixed CD/DP1 systems that identifies the source of ill-conditioning and prescribes reference-configuration row weights that restore a balanced Newton system. Second, it derives closed-form internal-force expressions and constitutive tangents for the SVK, Mooney–Rivlin, and Kelvin–Voigt models through a unified $\mathbf{P}(\mathbf{F})$ interface that is independent of element topology. Third, it assembles kinematics, constitutive response, external loading, frictional contact, and bilateral constraints into a single virtual-work framework, providing a consistent derivation of every force and reaction term that enters the equations of motion. Fourth, it recasts the backward-Euler constrained TL-FEA step as a velocity-level augmented-Lagrangian optimization problem: the stationarity conditions of a single scalar objective recover the discrete equations of motion, bilateral constraints are enforced through multiplier and penalty terms whose transpose action is assembled element-by-element without forming a dense global constraint matrix, and the structure of the resulting Newton system — including the positive semidefinite Gauss–Newton term and the indefinite DP1 curvature correction — is analyzed explicitly, linking the choice of linearization to solver requirements. Claims regarding computational efficiency, robustness, and large-scale numerical performance are addressed in the companion paper [18].
The remainder of the paper is organized as follows. Section 2 introduces the TL deformation map, the $\mathbf{N}(t)\,\mathbf{s}(\mathbf{u})$ interpolation, and the constraint primitives and engineering joints. Section 3 presents the stress measures and material models. Section 4 derives the virtual-work contributions for inertia, body forces, internal forces, concentrated loads, and surface tractions. Section 5 describes the frictional contact model. Section 6 covers numerical quadrature in the isoparametric setting. Section 7 formulates the augmented-Lagrangian time-discrete step. Section 8 summarizes the contributions. Appendices provide shape function data for the element types used in Part II, details of the joint constraint linearization, and closed-form tangent expressions for all material models.
2 Kinematics Aspects
2.1 The TL Deformation Map
Large-deformation multibody dynamics can be formulated in a continuum setting by describing, for each element, the deformation map $\mathbf{r}=\bm{\phi}(\mathbf{u};t)\in\mathbb{R}^{3}$ and the associated deformation gradient $\mathbf{F}(\mathbf{u};t)=\partial\mathbf{r}/\partial\mathbf{u}$, where $\mathbf{u}=[u,v,w]^{T}\in\mathbb{R}^{3}$ denotes a material point in the element reference configuration and $\mathbf{r}(\mathbf{u};t)$ gives its current location at time $t$.
Figure 1: Four configurations associated with one element in TL-FEA. The parent and reference configurations are auxiliary constructs used for interpolation and constitutive description; the initial and current configurations are the physical configurations at $t=0$ and at the current time, respectively.
Figure 1 distinguishes four configurations used throughout the paper. The reference configuration is the stress-free, strain-free configuration of the element, typically supplied by a mesh generator such as Gmsh [19] or TetGen [20]. The current configuration is described relative to a fixed inertial frame $Oxyz$ through the map $\mathbf{r}(u,v,w;t)=[x(u,v,w;t),y(u,v,w;t),z(u,v,w;t)]^{T}$. The initial configuration is the physical configuration at $t=0$, which may already carry stress. Finally, the parent (canonical) configuration is introduced for numerical integration; depending on the element, the parent-to-reference map is either isoparametric or affine.
In standard TL-FEA, the unknowns are the nodal quantities that parameterize the current configuration of the element. For a T10 tetrahedral element [14], these are the ten nodal positions $\mathbf{e}_{1}(t),\ldots,\mathbf{e}_{10}(t)\in\mathbb{R}^{3}$. Within flexible multibody dynamics, the Absolute Nodal Coordinate Formulation (ANCF) [21] is another TL-FEA instance, distinguished by the use of nodal positions together with spatial gradients as degrees of freedom. The ANCF literature spans beam, cable, shell, and solid-like elements, together with remedies for locking and related pathologies [22, 23, 24, 25, 26]. For present purposes, the relevant commonality is that both classical TL-FEA and ANCF describe a continuum map through the deformation gradient and strain measures referred to the reference configuration.
In the standard notation, the map $\bm{\phi}$ is approximated by interpolating the nodal unknowns as
$\mathbf{r}(\mathbf{u};t)=\mathbf{S}(\mathbf{u})\>\mathbf{e}(t)\;,$ (1)
where $\mathbf{S}(\mathbf{u})\in\mathbb{R}^{3\times 3n}$ is a block shape-function matrix and $\mathbf{e}(t)\in\mathbb{R}^{3n}$ collects the nodal degrees of freedom [5, 26]. For instance, for the T10 tetrahedron $n=10$, whereas for a fully parameterized two-node beam element $n=8$.
In this work, the same map is written instead as
$\mathbf{r}(\mathbf{u};t)=\mathbf{N}(t)\>\mathbf{s}(\mathbf{u})\;,$ (2)
where $\mathbf{N}(t)=[\mathbf{e}_{1}(t),\ldots,\mathbf{e}_{n}(t)]\in\mathbb{R}^{3\times n}$ stores the nodal unknowns columnwise and $\mathbf{s}(\mathbf{u})\in\mathbb{R}^{n}$ collects the scalar shape functions. In classical TL-FEA, the columns of $\mathbf{N}$ are nodal positions; in ANCF, they may include positions together with slope-related quantities. This representation is algebraically equivalent to Eq.˜1, but it will be used throughout because it leads to a compact expression for the deformation gradient, virtual displacements, and the Jacobian blocks that enter the constraint and internal-force calculations.
From Eq.˜2, the velocity, acceleration, and virtual displacement of the current point associated with $\mathbf{u}$ are
$\dot{\mathbf{r}}(u,v,w;t)=\dot{\mathbf{N}}(t)\>\mathbf{s}(u,v,w),\qquad\ddot{\mathbf{r}}(u,v,w;t)=\ddot{\mathbf{N}}(t)\>\mathbf{s}(u,v,w),$ (3a)
and
$\delta\mathbf{r}^{T}=\mathbf{s}^{T}(u,v,w)\,(\delta\mathbf{N})^{T}=\mathbf{s}^{T}(u,v,w)\begin{bmatrix}\delta{\mathbf{e}}_{1}^{T}\\ \delta{\mathbf{e}}_{2}^{T}\\ \vdots\\ \delta{\mathbf{e}}_{n}^{T}\end{bmatrix}\;$. (3b)
Let $s_{i}(u,v,w)$ denote the $i$-th entry of $\mathbf{s}(u,v,w)$ and define
$\mathbf{h}_{i}(u,v,w)\equiv\nabla_{\mathbf{u}}\,s_{i}(u,v,w)\in\mathbb{R}^{3},$ (4a)
and
$\mathbf{H}(u,v,w)=\begin{bmatrix}\mathbf{h}_{1}^{T}(u,v,w)\\ \mathbf{h}_{2}^{T}(u,v,w)\\ \vdots\\ \mathbf{h}_{n}^{T}(u,v,w)\end{bmatrix}=\frac{\partial\mathbf{s}(u,v,w)}{\partial\mathbf{u}}\in\mathbb{R}^{n\times 3}\;.$ (4b)
The deformation gradient then takes the form
$\mathbf{F}=\frac{\partial\mathbf{r}}{\partial\mathbf{u}}=\mathbf{N}(t)\>\mathbf{H}(u,v,w)=\sum_{i=1}^{n}\mathbf{e}_{i}(t)\,\mathbf{h}_{i}^{T}(u,v,w)\;.$ (5)
The strain measures used in the remainder of the paper are the right Cauchy–Green tensor and the Green–Lagrange strain tensor,
$\mathbf{C}\coloneqq\mathbf{F}^{T}\mathbf{F}\,$, (6a)
$\mathbf{E}\coloneqq\frac{1}{2}(\mathbf{C}-\mathbf{I})=\frac{1}{2}(\mathbf{F}^{T}\mathbf{F}-\mathbf{I})\,.$ (6b)
The expressions of $\mathbf{s}$ and $\mathbf{H}$ for several element types are provided in A.
2.2 Kinematic Constraints
We consider bilateral holonomic constraints built from four scalar geometric primitives—dot-product 1 (DP1), dot-product 2 (DP2), distance (DIST), and coordinate-difference (CD). These four primitives are sufficient to show how deformable-body joints are expressed in the proposed TL-FEA notation and how the corresponding Jacobian blocks follow from the point-evaluation operator Eq.˜2 and its first variation Eq.˜3b. Richer primitive libraries can be handled in the same way, but are not needed here.
Each primitive is posed as a scalar position-level condition $c(\mathbf{q},t)=0$. Since the discrete step is written as $\mathbf{q}_{n+1}=\mathbf{q}_{n}+h\,\mathbf{v}$, the nonlinear solve at time $t_{n+1}$ requires the constraint value $c(\mathbf{q}_{n+1},t_{n+1})$ and its Jacobian $J_{c}(\mathbf{q}_{n+1},t_{n+1})=\partial c/\partial\mathbf{q}$, which supplies the constraint-direction contributions entering the augmented-Lagrangian residual in Eq. (86). The purpose of this subsection is therefore threefold: to define the primitive constraints, to show how they are composed into standard engineering joints, and to make explicit the Jacobian accumulation and linearization issues that arise once dot-product constraints are present.
The primitives DP1, DP2, DIST, and CD are scalar compositions of point evaluations Eq.˜2 and point differences, so their Jacobians follow the same chain-rule pattern. For compactness, each primitive is stated through its defining scalar equation, its first variation, and the corresponding nonzero element-level Jacobian blocks.
2.2.1 Dot-product 1 (DP1)
DP1 enforces
Using Eq.˜3b at each point, the only nonzero Jacobian blocks are those associated with the four hosting elements. For example, the contribution associated with point $\mathcal{P}$ has the form
2.2.2 Dot-product 2 (DP2)
DP2 prescribes the dot product between a body-attached direction and a connector direction. In DP1, each direction is defined by two points that lie on the same body. In keeping with the nomenclature of [10], DP2 instead uses three points on one body and one point on the other. Let $\mathcal{P},\mathcal{Q},\mathcal{R}$ lie on body $b$ (elements $E,F,G$) and $\mathcal{T}$ lie on body $c$ (element $H$). Define
where $\mathbf{a}$ is a body-attached direction on body $b$ and $\mathbf{b}$ is a connector direction from point $\mathcal{R}$ on body $b$ to point $\mathcal{T}$ on body $c$. DP2 enforces
with first variation
2.2.3 Distance (DIST)
with $f(t)>0$. The first variation is
2.2.4 Coordinate-difference (CD)
The first variation is
so that the only nonzero Jacobian blocks are
2.2.5 Engineering Joints by Constraint Composition
Each scalar primitive contributes one scalar kinematic condition. Engineering joints are obtained by stacking such primitives. The degree-of-freedom counts reported in Table˜1 are local counts that hold when the assembled constraint rows are independent. In particular, the point sets used to define the joint must be regular: points used to define directions must remain distinct, the resulting direction vectors must be nonzero, off-axis directions must be linearly independent when required, and the assembled joint Jacobian must have full row rank. Under these assumptions, stacking $m$ independent scalar constraints removes $m$ relative degrees of freedom.
Table˜1 catalogs the primitive decomposition for six standard joint types.
Table 1: Primitive decomposition for engineering joints. CD denotes coordinate-difference (point-coincidence) constraints and DP denotes dot-product constraints. For the cylindrical and prismatic joints the notation a+ba{+}b indicates $a$ axis-parallelism (DP1-pattern) plus $b$ offset-collinearity (DP2-pattern) constraints.
Joint type CD DP Total $m$ Remaining DOF
Spherical 3 0 3 3 rotational
Universal 3 1 4 2 rotational
Revolute 3 2 5 1 rotational
Fixed (weld) 3 3 6 0
Cylindrical 0 2+22{+}2 4 1 rot. + 1 transl.
Prismatic 0 3+23{+}2 5 1 translational
The table should therefore be read as a catalog of constraint compositions, not as an unconditional guarantee. Degenerate point choices or near-singular configurations reduce the rank of the assembled Jacobian and change the effective mobility. We illustrate the construction with the revolute joint.
Revolute-joint point set.
Five material points define the joint. On body $b$: the attachment point $\mathcal{P}$ (element $E$) at the hinge location, and a nearby point $\mathcal{Q}$ (element $F$) such that the body-attached direction $\mathbf{a}=\mathbf{r}_{\mathcal{Q}}-\mathbf{r}_{\mathcal{P}}$ is aligned with the intended hinge axis in the reference configuration. This requires $\mathcal{P}$ and $\mathcal{Q}$ to be distinct, so that $\mathbf{a}\neq\mathbf{0}$. On body $c$: the attachment point $\mathcal{R}$ (element $G$) at the hinge location, and two off-axis points $\mathcal{S}$ and $\mathcal{T}$ (elements $H$ and $K$) defining directions $\mathbf{b}_{1}=\mathbf{r}_{\mathcal{S}}-\mathbf{r}_{\mathcal{R}}$ and $\mathbf{b}_{2}=\mathbf{r}_{\mathcal{T}}-\mathbf{r}_{\mathcal{R}}$. The directions $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ must be linearly independent and not both parallel to the hinge axis. Under these regularity conditions, the five scalar constraints introduced below are generically independent. The revolute-joint geometry is illustrated in Fig. 2.
Figure 2: Geometry of the revolute joint construction used in this subsection. Points $\mathcal{P},\mathcal{Q}$ on body $b$ define the body-attached axis direction $\mathbf{a}=\mathbf{r}_{\mathcal{Q}}-\mathbf{r}_{\mathcal{P}}$, while points $\mathcal{R},\mathcal{S},\mathcal{T}$ on body $c$ define the off-axis directions $\mathbf{b}_{1}=\mathbf{r}_{\mathcal{S}}-\mathbf{r}_{\mathcal{R}}$ and $\mathbf{b}_{2}=\mathbf{r}_{\mathcal{T}}-\mathbf{r}_{\mathcal{R}}$. Three CD constraints enforce point coincidence at the hinge location ($\mathbf{r}_{\mathcal{P}}=\mathbf{r}_{\mathcal{R}}$), and two DP1 constraints prescribe the relative orientation through $\mathbf{a}^{\mathsf{T}}\mathbf{b}_{1}=f_{1}$ and $\mathbf{a}^{\mathsf{T}}\mathbf{b}_{2}=f_{2}$, leaving one rotational degree of freedom about the hinge axis in the regular, full-rank case.
Constraint equations.
The revolute joint is defined by $m=5$ scalar constraints: three CD constraints for point coincidence and two DP1 constraints for relative orientation,
$\mathbf{c}^{\mathrm{rev}}=\begin{bmatrix}\mathbf{e}_{x}^{\mathsf{T}}(\mathbf{r}_{\mathcal{R}}-\mathbf{r}_{\mathcal{P}})\\[2.0pt] \mathbf{e}_{y}^{\mathsf{T}}(\mathbf{r}_{\mathcal{R}}-\mathbf{r}_{\mathcal{P}})\\[2.0pt] \mathbf{e}_{z}^{\mathsf{T}}(\mathbf{r}_{\mathcal{R}}-\mathbf{r}_{\mathcal{P}})\\[2.0pt] \mathbf{a}^{\mathsf{T}}\mathbf{b}_{1}-f_{1}\\[2.0pt] \mathbf{a}^{\mathsf{T}}\mathbf{b}_{2}-f_{2}\end{bmatrix}=\mathbf{0}\in\mathbb{R}^{5},$ (21)
where $f_{j}=\mathbf{a}_{0}^{\mathsf{T}}\mathbf{b}_{j,0}$ are constants precomputed from the reference configuration. Under the regularity conditions stated above, and provided the assembled joint Jacobian has local rank five, the three CD rows remove the three relative translations at the hinge, while the two DP1 rows remove two relative orientation components, leaving one rotational degree of freedom about $\mathbf{a}$. When the off-axis directions are chosen perpendicular to the axis (i.e., $f_{1}=f_{2}=0$), the DP1 constraints reduce to $\mathbf{a}^{\mathsf{T}}\mathbf{b}_{1}=0$ and $\mathbf{a}^{\mathsf{T}}\mathbf{b}_{2}=0$, forcing $\mathbf{a}\parallel(\mathbf{b}_{1}\times\mathbf{b}_{2})$.
Accumulated Jacobian blocks.
At the five constraint points, define the evaluation operators
$\bm{\Sigma}_{E}:=\mathbf{s}_{E}^{b}(\mathbf{u}_{\mathcal{P}})^{\mathsf{T}}\otimes\mathbf{I}_{3},\qquad\bm{\Sigma}_{F}:=\mathbf{s}_{F}^{b}(\mathbf{u}_{\mathcal{Q}})^{\mathsf{T}}\otimes\mathbf{I}_{3},\qquad\bm{\Sigma}_{G}:=\mathbf{s}_{G}^{c}(\mathbf{u}_{\mathcal{R}})^{\mathsf{T}}\otimes\mathbf{I}_{3},\qquad\bm{\Sigma}_{H}:=\mathbf{s}_{H}^{c}(\mathbf{u}_{\mathcal{S}})^{\mathsf{T}}\otimes\mathbf{I}_{3},\qquad\bm{\Sigma}_{K}:=\mathbf{s}_{K}^{c}(\mathbf{u}_{\mathcal{T}})^{\mathsf{T}}\otimes\mathbf{I}_{3}.$ (22)
Since $\mathcal{P}$ participates in both the CD and DP1 rows, and $\mathcal{R}$ does likewise, the corresponding hosting elements receive accumulated $5\times 3n_{E}$ and $5\times 3n_{G}$ Jacobian blocks:
$\frac{\partial\mathbf{c}^{\mathrm{rev}}}{\partial\mathbf{e}_{E}^{b}}=-\begin{bmatrix}\mathbf{e}_{x}^{\mathsf{T}}\\[2.0pt] \mathbf{e}_{y}^{\mathsf{T}}\\[2.0pt] \mathbf{e}_{z}^{\mathsf{T}}\\[2.0pt] \mathbf{b}_{1}^{\mathsf{T}}\\[2.0pt] \mathbf{b}_{2}^{\mathsf{T}}\end{bmatrix}\bm{\Sigma}_{E},\qquad\frac{\partial\mathbf{c}^{\mathrm{rev}}}{\partial\mathbf{e}_{G}^{c}}=\begin{bmatrix}+\mathbf{e}_{x}^{\mathsf{T}}\\[2.0pt] +\mathbf{e}_{y}^{\mathsf{T}}\\[2.0pt] +\mathbf{e}_{z}^{\mathsf{T}}\\[2.0pt] -\mathbf{a}^{\mathsf{T}}\\[2.0pt] -\mathbf{a}^{\mathsf{T}}\end{bmatrix}\bm{\Sigma}_{G}.$ (23)
The remaining elements contribute only through the DP1 rows:
$\frac{\partial\mathbf{c}^{\mathrm{rev}}}{\partial\mathbf{e}_{F}^{b}}=\begin{bmatrix}\mathbf{0}^{\mathsf{T}}\\[1.0pt] \mathbf{0}^{\mathsf{T}}\\[1.0pt] \mathbf{0}^{\mathsf{T}}\\[1.0pt] \mathbf{b}_{1}^{\mathsf{T}}\\[1.0pt] \mathbf{b}_{2}^{\mathsf{T}}\end{bmatrix}\bm{\Sigma}_{F},\qquad\frac{\partial\mathbf{c}^{\mathrm{rev}}}{\partial\mathbf{e}_{H}^{c}}=\begin{bmatrix}\mathbf{0}^{\mathsf{T}}\\[1.0pt] \mathbf{0}^{\mathsf{T}}\\[1.0pt] \mathbf{0}^{\mathsf{T}}\\[1.0pt] \mathbf{a}^{\mathsf{T}}\\[1.0pt] \mathbf{0}^{\mathsf{T}}\end{bmatrix}\bm{\Sigma}_{H},\qquad\frac{\partial\mathbf{c}^{\mathrm{rev}}}{\partial\mathbf{e}_{K}^{c}}=\begin{bmatrix}\mathbf{0}^{\mathsf{T}}\\[1.0pt] \mathbf{0}^{\mathsf{T}}\\[1.0pt] \mathbf{0}^{\mathsf{T}}\\[1.0pt] \mathbf{0}^{\mathsf{T}}\\[1.0pt] \mathbf{a}^{\mathsf{T}}\end{bmatrix}\bm{\Sigma}_{K}.$ (24)
These expressions make explicit how the primitive-level Jacobian rows combine into a composite-joint Jacobian. In particular, the algebraic coupling at the shared attachment points is already present at the element-block level; the global joint Jacobian is obtained by scattering these local blocks into the global row structure.
Constraint row normalization.
A mixed CD/DP1 constraint set introduces a scale mismatch. CD constraint values are $\mathcal{O}(\|\Delta\mathbf{r}\|)$, whereas DP1 values scale as $\mathcal{O}(\delta^{2})$ when the direction vectors are defined by material fibers of length $\delta$, typically chosen as a small fraction of the element diameter. The corresponding Jacobian entries scale as $\mathcal{O}(1)$ for CD and $\mathcal{O}(\delta)$ for DP1. Without correction, the translational rows dominate both the residual and the Gauss–Newton term, and the angular part of the Newton step becomes poorly conditioned. We therefore apply constant row weights computed once from the reference configuration:
$w_{k}=1\qquad(k=1,2,3;\;\text{CD}),\qquad w_{3+j}=\frac{1}{\sqrt{|\mathbf{a}_{0}|^{2}+|\mathbf{b}_{j,0}|^{2}}}\qquad(j=1,2;\;\text{DP1}).$ (25)
With $|\mathbf{a}_{0}|=|\mathbf{b}_{j,0}|=\delta$, one obtains $w_{3+j}\approx 1/\delta$. If $W_{\mathrm{rev}}=\mathrm{diag}(w_{1},\ldots,w_{5})$, the row-normalized revolute constraint vector and element Jacobian blocks are
$\widehat{\mathbf{c}}^{\mathrm{rev}}=W_{\mathrm{rev}}\,\mathbf{c}^{\mathrm{rev}},\qquad\widehat{J}^{\mathrm{rev}}_{X}=W_{\mathrm{rev}}\,\frac{\partial\mathbf{c}^{\mathrm{rev}}}{\partial\mathbf{e}_{X}},\qquad X\in\{E,F,G,H,K\}.$ (26)
This normalization balances the translational and angular rows at the level of both residual and Jacobian. A more detailed scaling argument is given in B.
Residual insertion and consistent linearization.
The normalized blocks in Eq.˜26 enter the augmented-Lagrangian residual through an element-wise scatter. If $\widehat{\bm{\eta}}^{\mathrm{rev}}=\widehat{\bm{\lambda}}^{\mathrm{rev}}+\rho\,\widehat{\mathbf{c}}^{\mathrm{rev}}$, each hosting element receives
$\mathbf{g}_{X}\;\mathrel{+}=\;h\,\bigl(\widehat{J}^{\mathrm{rev}}_{X}\bigr)^{\mathsf{T}}\widehat{\bm{\eta}}^{\mathrm{rev}},\qquad X\in\{E,F,G,H,K\}.$ (27)
Accordingly, the global transpose action $\mathbf{C}_{q}^{\mathsf{T}}\widehat{\bm{\eta}}$ is never formed as a dense matrix; it is assembled element-by-element in exactly the same way as internal and external force contributions.
The Newton system requires the derivative of the constraint contribution with respect to $\mathbf{v}$. With $\mathbf{q}=\mathbf{q}_{n}+h\,\mathbf{v}$ and with $\mathbf{c}$ and $\mathbf{C}_{q}$ denoting the row-normalized constraints and Jacobian used in the solve, one obtains
$\frac{\partial\mathbf{g}_{\mathrm{con}}}{\partial\mathbf{v}}=\underbrace{h^{2}\rho\,\mathbf{C}_{q}^{\mathsf{T}}\mathbf{C}_{q}}_{\displaystyle\mathbf{H}_{\mathrm{GN}}}+\underbrace{h^{2}\sum_{i=1}^{m}\eta_{i}\,\nabla_{q}^{2}c_{i}}_{\displaystyle\mathbf{H}_{\mathrm{curv}}},$ (28)
where $\eta_{i}=\lambda_{i}+\rho\,c_{i}$. Each CD constraint is linear in $\mathbf{q}$, so its Hessian vanishes. A DP1 constraint is quadratic in $\mathbf{q}$, and its Hessian is therefore constant, sparse, and in general indefinite. The explicit block form is given in B. The Gauss–Newton contribution $\mathbf{H}_{\mathrm{GN}}$ is positive semidefinite. The curvature correction $\mathbf{H}_{\mathrm{curv}}$ is the term required for a consistent Newton linearization, but because it is indefinite, the full Newton matrix
$\mathbf{H}=\frac{1}{h}\mathbf{M}+h\,\mathbf{K}_{t}+h^{2}\rho\,\mathbf{C}_{q}^{\mathsf{T}}\mathbf{C}_{q}+h^{2}\textstyle\sum_{i=1}^{m}\eta_{i}\,\nabla_{q}^{2}c_{i}$ (29)
[…]